# Accelerating Speculative Decoding using Dynamic Speculation Length

### v1

@@ Appendix A Oracle SL Analysis
Fig. 6 shows bar charts of the average oracle SL over the normalized index of the Speculation Iteration. We calculate the bars as follows. For each prompt of a dataset, we have its corresponding sequence of oracle SLs. The length of the sequence is equal to the number of Speculation Iterations. For example, consider a prompt with oracle SLs $\langle 7,3,13,21,8\rangle$. Its normalized index is $\langle 0,0.25,0.5,0.75,1\rangle$. The bars represent the average oracle SL of buckets of size $0.0001$.
@@ Appendix D Classifier
D.1 Classifier Training Data
In order to train the model we created training data using the following method: the target model followed the ‘standard’ autoregressive approach to generate tokens based on the input prompt, whereas the draft model iteratively generated a single token per iteration. In each iteration, the draft model generates the next token based on the concatenation of the input prompt and the tokens subsequently generated by the target model. Draft tokens resembling the target token at the same position were labeled ‘accepted’, while others were labeled ‘rejected’. Corresponding features were extracted for both draft and target tokens, encompassing the top-k probabilities of the vocabulary distribution, the entropy associated with these probabilities, and the token position value counted from the beginning of the generation process.
@@ Appendix E Additional Implementation Details
For every dataset, DISCO classifier is trained on the train set; threshold $\tau$ and $SL_{max}$ hyper-parameters are fine-tuned on the validation set optimizing the latency. Optimal static SL is estimated on the validation set. Latency results are reported on the test set.

### v4

@@ Appendix A Oracle SL Analysis
We hypothesized that later tokens are more predictive but eventually found only a relatively weak correlation, as Fig. 2 and Fig. 5 show. The figures are bar charts of the average oracle SL over the normalized index of the Speculation Iteration. We calculate the bars as follows. For each prompt of a dataset, we have its corresponding sequence of oracle SLs. The length of the sequence is equal to the number of Speculation Iterations. For example, consider a prompt with oracle SLs $\langle 7,3,13,21,8\rangle$. Its normalized index is $\langle 0,0.25,0.5,0.75,1\rangle$. The bars represent the average oracle SL of buckets of size $0.0001$.
@@ Appendix D Classifier
D.1 Feature Extraction from Training Data for Classifier Training
To train the model for each dataset, MBPP, CNN-DM, and Alpaca, we used the corresponding training set of each dataset. For each token in each training set, we extracted a boolean label (accepted/rejected) and a list of features in the following manner: We ran the target model using the standard autoregressive approach for generating tokens based on the input prompt. In contrast, the draft model iteratively generated only a single token per iteration. During each iteration, the draft model generated this token based on the concatenation of the input prompt and the tokens subsequently generated by the target model. Draft tokens that resembled the target token at the same position were labeled "accepted," while others were labeled "rejected." Corresponding features were extracted for both draft and target tokens, encompassing the top-k probabilities of the vocabulary distribution, the entropy associated with these probabilities, and the token position value counted from the beginning of the generation process.
@@ Appendix E Additional Implementation Details
For every dataset, DISCO classifier is trained on the train set; threshold $\tau$ and $\text{SL}_{\max}$ hyper-parameters are fine-tuned on the validation set optimizing the latency. Optimal static SL is estimated on the validation set. Latency results are reported on the test set.